OpenStudy (anonymous):

maximize the objective function p = 3x+y sibject to the following constraints 4x+5y≤9 2x-y≥1 x≥0 y≥0 on my teachers answer sheet ht has 1/2,0 as his corner points 2.25,0 as another 1,1(which is the intersection of the two lines), how did my teacher get the other two corner points?

OpenStudy (anonymous):

by the simultaneous solution of both lines

OpenStudy (anonymous):

i saw the 1,1 with the intersection but did not find the (2.25,0) and the (1/2,)

OpenStudy (anonymous):

1/2,0 i mean

OpenStudy (anonymous):

both the points lie on x-axis

OpenStudy (anonymous):

mean they are the intersection of lines with x-axis

OpenStudy (anonymous):

i got the 1,1 intersection using my calculator but the 1/2,0 and 2.25, 0 did not come up. how would i find those?

OpenStudy (anonymous):

\[4x+5y \le 9\]

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

y=0

OpenStudy (amistre64):

2x-y≥1 is this a typo?

OpenStudy (anonymous):

\[2x-y \ge 1\]

OpenStudy (anonymous):

thats what my teacher put

OpenStudy (anonymous):

oh i have to write the equtions instead of inequalities

OpenStudy (anonymous):

the answers are y ≤ 9-4x/5 and y ≤ 2x-1

OpenStudy (anonymous):

4x+5y=9 y=0 gives x=9/4=2.25 point of intersection of the line and x-axis is (2.25,0)

OpenStudy (anonymous):

likewise for the second eqn 2x-y=1 y=0 x=1/2 point is (1/2,0)

OpenStudy (anonymous):

is the 5 in the deonominator a typo on his part?

OpenStudy (anonymous):

5 in denominator?

OpenStudy (anonymous):

this is how we get the corner points

OpenStudy (anonymous):

on his answer key it says 9-4x/5

OpenStudy (anonymous):

that makes more sense thanks a lot

OpenStudy (anonymous):

so should i change the answer on the key where it says 9-4x/5?

OpenStudy (anonymous):

is it y=9-4x/5?

OpenStudy (anonymous):

it was originally 4x+5y≤ 9

OpenStudy (anonymous):

fine..what next?

OpenStudy (anonymous):

to draw the line transform it into eqn 4x+5y=9

OpenStudy (anonymous):

then he has his answer as y ≤ 9-4x/5

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

what next?

OpenStudy (anonymous):

the x should then be move to the other side

OpenStudy (anonymous):

the 4x

OpenStudy (anonymous):

which x?actually m not getting you what you actually want

OpenStudy (anonymous):

u are to maximize the objective function subject to some constraints...right?

OpenStudy (anonymous):

thats correct

OpenStudy (anonymous):

for that u r given two lines and first quadrant

OpenStudy (anonymous):

we draw the lines n mark the feasible region, just like explained in ur previous post

OpenStudy (anonymous):

maximum value liea at one of the corner points

OpenStudy (anonymous):

lies*

OpenStudy (anonymous):

ok that makes more sense

OpenStudy (anonymous):

so r u fine with it now?